question:
to multiiply 2 1/2 X 3 1/4
can you multiply 2x3 and 1/2 by 1/4
why does this not work

Answer by liz_gcmama(8)

to multiiply 2 1/2 X 3 1/4
can you multiply 2x3 and 1/2 by 1/4
The short answer is no.
To solve this problem you need 2 1/2 times the 3 and 2 1/2 times the 1/4. The easiest way to solve this problem is to make them improper fractions.
2 1/2 = 2+ 1/2 = 4/2 + 1/2 = 5/2
3 1/4 = 3 + 1/4 = 12/4 + 1/4 = 13/4
So the problem now looks like
5/2 X 13/4
now you can multiply 5x13 to get 65
then you multiply 2x4 to get 8
so you end up with 65/8.
You can simplify this further. We know that 8*8 is 64
65/8=64/8+1/8 = 8+1/8
So the answer is 8 1/8